# A Computational Intelligence Approach to Personal Scoring, Risk Stratification and Immune Age Characterisation from Common Blood Markers

### v1

@@ Abstract
We introduce a family of indices based on a distance metric and closed learning algorithm for detecting rapid, small (precise) and personalised changes in blood test results. The indexes have the advantage of being fully explainable, unlike those featured in other statistical or machine-learning approaches. This metric takes the form of a quantitative measure or score that aggregates data from a set of numeric blood test results, combining and condensing them into a single numerical value that allows rapid triage, sorting, and immune age assessment. The main purpose of the score is to learn from, monitor, and provide a quick assessment of divergence from population and personalised reference values. We demonstrate the efficacy of the score against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from the CDC NHANES survey spanning 13 years, from 2003 to 2016. We find that the score we introduce is informative when it comes to distinguishing healthy individuals from those with diseases, both self-reported and as manifest in abnormal blood test results, thereby providing an entry-level score for patient triaging. We show that the score varies over time and is correlated with biological age, leading to definition of an immune age as the inverse function of this relationship when taking different baskets of analytes based only on CBC test results, affording clinical evidence of its potential relevance to personalised precision medicine and individualised predictive healthcare.
Keywords: personalised medicine, quantitative precision medicine, pro-active healthcare, early detection, blood risk score, risk stratification, individual immune system assessment, immune age
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores for patient sorting and triaging are often qualitative, and are needed for the selection of patient pathways and health monitoring choices across all levels of health care.
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a Complete/Full Blood Count (CBC/FBC) risk score shown to be associated with bleeding and in-hospital mortality based on variations in red blood cell size, such variations, originally used as an anaemia predictor [26], being often reported in FBC/CBC panels. FBC/CBCs have been found to contain enough information to yield clinically useful scores [3, 18]. At the same time, it has also been shown that although a CBC is a low-cost laboratory test that is almost universally used, its risk predictive information content is often underused [29].
Regular blood testing has been associated with short term mortality and with the ability to significantly improve emergency department triage [22]. Most risk scores depend on multiple factors external to a single test result in and of itself, and have proven useful when available, even in the context of the immune system [34]. A number of these scores are based on such factors as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. While numerical scores and colour-based scores are relatively common, to the authors’ knowledge no single general numerical and colour-based blood test risk assessment score exists that is based on a combination of distances from population or personalised average values, in particular for CBC and immune-related data.
In medical practice, a simple risk score based solely on blood test results and agnostic in regards to specific condition while still being able to quantify the results of a blood test, has not been proposed, to the authors’ knowledge. Here we introduce a general numerical and colour-based risk stratification score to quantify any number of abnormal blood markers. The risk model introduced is not data expensive when run ab initio, and yet it is shown to be informative, and its implementation is easily interpretable, while allowing for adaptation to enhance sophistication.
@@ 2 Methods
We call the risk assessment score introduced here the ”immune score”, given its connection to the immune system by way of the counts of blood cell subpopulations, in particular as registered in the most popular blood test, the Full Blood Count (FBC) or Complete Blood Count (CBC).
The immune score constitutes a dimensional reduction technique based on a single real-value number and a colour scheme that takes a multidimensional blood test space with dimension size equal to the number of analytes, where each value of a marker or analyte corresponds to the coordinate of that value in that dimension. The score itself can be seen as the norm of a suitable transformation of a vector pinpointing the health status of a patient in that space, relative to the test and set of markers or analytes. So the numerical vector value integrates all the analyte’s space dimensions.
@@ 4.2 Adaptive reference values
5 Immune age
The immune system plays an important role in protecting against infections and in the maintenance of human health. Blood is the medium in which most immune cells circulate through the body.
Ageing is associated with complex changes in and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described [24].
Worldwide, as lifespans are generally increasing, there is a concomitant increase in age-related morbidities, particularly cardiovascular diseases, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune functions.
@@ 5 Immune age
5.1 Immune Score Variation Over time
@@
To test the hypothesis, we selected two cohorts comprising all the entries with ages ranging between 20 to 30 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
@@
5.2 Immune age based on immune score
@@
As shown in Figure 14, prediction error rates are much lower in the healthy NHANES data, in comparison to the unhealthy groups, for both sexes, and can be therefore used to quantify healthy patient ages to some extent. The prediction error rates for the immune-score age, was overestimated roughly within 7 years, for the healthy groups, in both sexes. Thus, the error rate of age prediction for healthy people is small, with a mean error rate of 7.33±1.21 for females, and 7.62 ±1.17 for healthy males. Further, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. The Pearson correlation statistics on the predicted age error rates, by RBC count health status, as discussed before, was an R-square value of 0.293, and 0.168, for males (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and non-significant when the self-reported health measures were used to separate the health status groups but were significant when taking abnormal CBC values only.
@@ 6 Conclusion
We have introduced risk-assessment scores and studied their behaviour against typical synthetic and empirical disease cases. We have also introduced a learning procedure to adapt normal versus abnormal reference values to personalised ranges for more precise and individual medical assessment and demonstrated the capabilities and applications of the scores also for triaging purposes against a popular public database showing how the scores separate healthy from unhealthy people even in the face of significant noise coming from self-reported health status. The numerical score was also shown to be able to define an ‘immune age’, with some significant differences for males and females who are known to have greater fluctuations related to e.g. menopause.
The approaches were shown to be informative for quick and entry-level patient health sorting and triaging purposes. In the case of the immune age, the age of healthy people was very closely predicted mostly by RBC as a component of the CBC, and for unhealthy both by reported or with actual abnormal CBC values, the score was not predictive hence indicating that unhealthy people will find a mismatch between their score age prediction and their actual biological age.
Unlike statistical or traditional machine learning approaches, these scores are transparent so clinicians and health professionals, as well as health consumers, can fully understand how they mechanistically work and allow calculation by hand with each step justified by a mathematical interpretation that can be assessed in a clinical context for rapid screening and decision support.

### v2

# An Adaptive Computational Intelligence Approach to Personalised Health Assessment and Immune Age Characterisation from Common Haematological Markers

@@ Abstract
We introduce an approach to creating a simulated digital twin of the health of an individual based on a learning adaptive algorithm that learns the optimal reference values of the blood panel over time and associates an immune age score to compare with where if the biological age is lower than the score, an indication of wellness is provided. The score may also be useful for classification purposes. We demonstrate its efficacy against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from 100K data records in the CDC NHANES survey that spans 13 years, from 2003 to 2016. We find that the score we introduce is informative when distinguishing healthy individuals from those with diseases, both self-reported and abnormal blood tests manifested, providing an entry-level score for patient triaging. We show that the score varies over time and is correlated with biological age, leading to the definition of an immune age as the inverse function of this relationship when different sets of analytes are taken only based on the results of the CBC test, providing clinical evidence of its potential relevance to the results of the CBC test, providing precision medicine and for personalised predictive healthcare.
Keywords: personalized health, predictive medicine, twin digital health, blood risk score, risk stratification, individual assessment, immune age
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores for patient sorting and triaging are often qualitative and are needed to select patient pathways and health monitoring choices across all levels of health care.
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a complete / complete blood count (CBC / FBC) risk score shown to be associated with bleeding and in-hospital mortality based on red blood cell size, originally used as an anaemia predictor [26], and is often reported in FBC/CBC panels.
FBC/CBCs have been found to contain enough information to produce clinically useful scores [3, 18]. At the same time, it has also been shown that although a CBC is a low-cost laboratory test that is almost universally used, its risk predictive information content is often underused [29].
Regular blood testing has been associated with short-term mortality and with the ability to significantly improve emergency department triage [22]. Most risk scores depend on multiple factors external to a single test result in itself and have proven useful when available, even in the context of the immune system [34]. A number of these scores are based on such factors as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. While numerical scores and colour-based scores are relatively common, to the authors’ knowledge no single general numerical and colour-based blood test risk assessment score exists that is based on a combination of distances from population or personalised average values, in particular for CBC and immune-related data.
In medical practice, a simple risk score based solely on blood test results and agnostic with respect to a specific condition while still being able to quantify the results of a blood test has not been proposed, to the authors’ knowledge. Here we introduce a general numerical and colour-based risk stratification score to quantify any number of abnormal blood markers. The risk model introduced is not data expensive when run ab initio, and yet it is shown to be informative, and its implementation is easily interpretable, while allowing for adaptation to enhance sophistication.
@@ 2 Methods
Here we introduce a family of indices based on a distance metric and closed learning algorithm for detecting rapid, small (precise), and personalised changes in blood test results. The indexes have the advantage of being fully explainable, unlike those featured in other statistical or machine-learning approaches. This metric takes the form of a quantitative measure or score that aggregates data from a set of numeric blood test results, combining and condensing them into a single numerical value that allows rapid triage, sorting, and immune age assessment. The main purpose of the score is to learn from, monitor, and provide a quick assessment of divergence from population and personalised reference values.
We will call this risk assessment score ”immune score”, given its connection to the immune system through blood cell subpopulation counts, in particular as registered in the most popular blood test, the Full Blood Count (FBC) or the Complete Blood Count (CBC).
The immune score constitutes a dimensional reduction technique based on a single real-value number and a colour scheme that takes a multidimensional blood test space with dimension size equal to the number of analytes, where each value of a marker or analyte corresponds to the coordinate of that value in that dimension. The score itself can be seen as the norm of a suitable vector transformation that pinpoints the health status of the patient in that space, relative to the test and the set of markers or analytes. Thus, the numerical vector value integrates all the analyte’s space dimensions.
@@ 4.2 Adaptive reference values
5 Immune age estimation
The immune system plays an important role in protecting against infections and in maintaining human health. Blood is the medium in which most immune cells circulate through the body.
Ageing is associated with complex changes and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described before[24].
As lifespan increases in many countries, there is a concomitant increase in age-related morbidities, particularly cardiovascular diseases, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune functions.
@@ 5 Immune age estimation
5.1 Immune Score Over time
@@
To test the hypothesis, we selected two cohorts comprising all the entries with ages ranging between 20 to 30 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
@@
5.2 Immune age based on the immune score
@@
As shown in Figure 14, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes and therefore can be used to quantify the ages of healthy patients to some extent. Prediction error rates for the immune score age were overestimated by approximately 7 years, for the healthy groups, in both sexes. Thus, the age prediction error rate for healthy people is small, with a mean error rate of 7.33 $\pm$ 1.21 for females, and 7.62 $\pm$ 1.17 for healthy males. Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. Pearson’s correlation statistics on the age error rates predicted, by the health status of the red blood cell count, as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
Figure 12: Greatest predictors and separators of age for healthy and unhealthy populations. The abrupt colour change is an artifact of typical normal versus abnormal values. Those in transition through the typical normal interval are missed and only recovered on one side or the other once they have already transitioned. These plots show that these dynamics are strongly predictive of age and health.
Table 5: Statistics of each CBC analyte as a predictor or discriminant of health and age
Counts Healthy Unhealthy Age
BASO 0.0051 0.0166 0.0004
EOS 0.0012 2.00E-05 0.0049
HCT 0.0036 0.0055 0.0696
HGB 0.1179 0.1397 0.0635
LYMP 0.0429 0.4983 0.0928
MCH 0.0831 0.0001 0.1538
MCHC 0.0025 0.0043 0.0047
MCV 0.2661 0.337 0.2186
MONO 0.0121 0.0004 0.0073
NEUT 0.0084 0.0069 0.018
RBC 0.0268 1.00E-05 0.0009
PLT 0.0921 0.3334 0.1029
Table 6: Best predictors for each category
Healthy Unhealthy Age
MCV LYMP MCV
HGB MCV MCH
PLT PLT PLT
MCH HGB LYMP
LYMP BASO HCT
RBC NEUT HGB
MONO HCT NEUT
NEUT MCHC MONO
BASO MONO EOS
HCT MCH MCHC
MCHC EOS RBC
EOS RBC BASO
@@ 6 Conclusion
We have introduced risk-assessment scores and studied their behaviour against typical synthetic and empirical disease cases. We have introduced a learning procedure to adapt normal versus abnormal reference values to personalised ranges for more precise and individual medical assessment and demonstrated the capabilities and applications of the scores for triaging purposes against a popular public database showing how the scores separate healthy from unhealthy cohorts both self-reported with significant noise and determined by typical normal reference values. The approaches were shown to be informative for quick and entry-level patient health sorting and triaging.
The numerical score was shown to be able to define an ‘immune age’, with some differences for males and females who are known to show greater fluctuations due to e.g. menopause. In the case of estimation of immune age, the biological age of healthy people was closely predicted by common hematological markers typically measured in FBC or CBC, and for unhealthy self-reported or by typical abnormal reference values, the score was not predictive of biological age, comforming with the theoretical expectation of a good predictor of health by divergence between estimated immune and actual biological age.
Unlike statistical or traditional machine learning and black-box approaches, the approach introduced is transparent to clinicians, health professionals, and health consumers who can follow the calculations and interpretation to be assessed in a clinical context for rapid screening and decision support.